OpenStudy (anonymous):

What is the sum of the arithmetic sequence 8, 15, 22 …, if there are 26 terms?

OpenStudy (anonymous):

First term is 8 here.. Common difference is 15 - 8 = 7 Use: \[\large Sum = \frac{n}{2}(2a + (n-1)d)\] a = 8 d= 7 and n= 26 can you do it??

OpenStudy (anonymous):

Trying or not @RoLexx

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Where you do not get it??

OpenStudy (anonymous):

You have all the values just plug in the formula..

OpenStudy (anonymous):

or you think this is the formula: \[\large Sum = \frac{n(first \quad + \quad last)}{2}\]

OpenStudy (anonymous):

Both the formulas are same..

OpenStudy (anonymous):

Can you plug in the values in the very first formula: a = 8 d = 7 and n = 26 Find the Sum..

OpenStudy (anonymous):

See: \[\large Sum = \frac{26}{2}(2(8) + (26-1)(7))\] \[Sum = 13(16 + 25 \cdot 7) \implies 13(16 + 175)\]

OpenStudy (anonymous):

Use calculator and fine the value of : \(13 \times 191\) = ??

OpenStudy (anonymous):

2483

OpenStudy (anonymous):

Absolutely.. Well Done..

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Welcome dear..